I'm just simply stuck. I can think through many parts of it in my mind. I can't remember how to put it in a formula.
How many ounces of 5% alcohol solution must be mixed with 16 ounces of a 10% alcohol solution to make a 6% alcohol solution?

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

How many ounces of 5% alcohol solution must be mixed with 16 ounces of a 10% alcohol solution to make a 6% alcohol solution?
--------------------------
Equation:
alcohol + alcohol = alcohol
0.05x + 0.10*16 = 0.06(x+16)
Multiply thru by 100 to get:
5x + 10*16 = 6x + 6*16
x = 4*16
x = 64 ounces (amt. of 5% solution in the mixture)
=======================
Cheers,
Stan H.

Answer by richwmiller(17219)


let x be the amount in ounces of 5% solution
the total solution will be 16 ounces plus x
.05x+.10(16)=.06*(x+16)
You could multiply both sides by 100 to get rid of the decimals